OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $14,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $2,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain

OpenStudy (anonymous):

option 1 i think its 60 dollars monthly if he leaves 2,000 down payment so 12,000/6 =200

OpenStudy (anonymous):

divided by 60 not 6

OpenStudy (anonymous):

Option one: pay 2000 down. Now you owe 14000-2000= 12000 to get the monthly payment, divide 12000/60 Option two: no down payment. So you owe 14000. Interest is 2% per year for 6 years: P*r*t=14000*.02*6= total interest that will be charged. Add interest to 14000 to get total amount of loan To get the monthly payment, divide by 72. Hoping this helps! (I got 200 for option 1, and 217.78 for option2)